Approximately how much greater in mass is a D2O molecule than an H2O molecule? (Assume the oxygen atoms have a mass number of 16. Deuterium, D, is an isotope of hydrogen with a mass number of 2, 2H.)

in amu
Answer #1

mass of D = 2

mass of O = 16

mass of D2O = 2*2 +16 = 20g/mole

mass of H2O = 1*2 + 16 = 18g/mole

mass of D2O is 2g/mole greater than mass of H2O

answer is 2 greater than H2O